The following table shows the actual demand observed over the last 11 years Yea Demand 7 9 5 10 13 7 13 14 9 11 8 This exercise contains only parts c, and d b) Using the 3-year moving average, provide the forecast from periods 4 through 12 (ound your responses to one decimal place. Year Forecast ear weighted moving average with weights 0.20, 0.35, and 0.45, using 0.45 for the most recent period, provide the forecast from periods 4 through 12 (round your responses to two decimal places c) Using the 3-y Yea Forecast d) Mean absolute deviation for the forecast developed using the 3-year moving average is 2.86 and for the 3y weighted moving average is 3.08. Based on this information, the better forecast is achieved using the ear approach.

The following table shows the actual demand observed over the last 11 years 

Year1234567891011
Demand7951013713149118
This exercise contains only parts b, c, and d.

b) Using the 3-year moving average, provide the forecast from periods 4 through 12 (round your responses to one decimal place). 

Year456789101112
Forecast








c) Using the 3-year weighted moving average with weights 0.20. 0.35. and 0.45. using 0.45 for the most recent period provide the forecast from periods 4 through 12.

Year456789101112
Forecast








d) Mean absolute deviation for the forecast developed using the 3-year moving average is 2 86 and for the 3-year weighted moving average is 3.08. Based on this information the better forecast is achieved using the approach
Answer #1

b) Using three year moving average,forecast from periods 4 through 12 are :

  • Period 4 = (7+9+5) / 3 = 7
  • Period 5 = (9+5+10) / 3 = 8
  • Period 6 = (5+10+13) / 3 =9.3
  • Period 7 = (10+13+7) / 3 = 10
  • Period 8 = (13+7+13) / 3 = 11
  • Period 9 = (7+13+14) / 3 = 11.3
  • Period 10 = (13+14+9) / 3 = 12
  • Period 11 = (14+9+11) / 3 =11.3
  • Period 12 = (9+11+8) / 3 = 9.3

c) Using 3-year weighted moving average the forecast from period 4 through 12 are:

  • Period 4 = (7X0.20) + (9X0.35) + (5X0.45) = 6.8
  • Period 5 = (9X0.20) + (5X0.35) + (10X0.45) =8.05
  • Period 6 = (5X0.20) + (10x0.35) + (13X0.45) =10.35
  • Period 7 = (10X0.20) + (13X0.35) + ( 7X0.45) = 9.7
  • Period 8 = (13X0.20) + (7X0.35) + (13X0.45) =10.9
  • Period 9 = (7X0.20) + (13x0.35) + (14x0.45) =12.25
  • Period 10 = ( 13X0.20) + (14X0.35) + (9X 0.45) =11.55
  • Period 11 = (14X0.20) + (9X0.35) + (11X0.45) = 10.9
  • Period 12 = (9X0.20) + (11X0.35) + (8X0.45) = 9.25

d) Based on the given information the better forecast is achieved using the three year moving average approach as it yields the lower Mean Absolute Deviation.